solve by factoring
x^2-10x=-16

Answer by praseenakos@yahoo.com(507)

QUESTION:
solve by factoring
x^2-10x=-16
ANSWER:
x^2-10x=-16

Add +16 on both sides of the equation

==> x^2-10x + 1`6= -16 + 16

==> x^2-10x + 16 = 0

This is a quadratic equation.

We can use quadratic fromula as well as factorisation method.

Factorisation method==>

x^2-10x + 16 = 0

First of all split the middle term.

For that find out two numbers whose sum is - 10 ( coefficient of x ) and whose product is +16 ( constant term )


Such two numbers are -8 and -2(( To get these numbers consider the factors of +16 , constant terms and try with different sign untill you get the numbers.)


Because,

-8 * -2 = -16 and -8 + ( -2 ) = -10

Now rewrite the given equation as follows,

x^2 - 8x - 2x + 16 = 0

Now group the terms as follows,


(x^2 - 8x ) - (2x - 16 ) = 0 ( when we put negative sign before the brackets, the sign inside the brackets will change)

Now take out common factor from each brackets.

==> x ( x - 8 ) - 2 ( x - 8 ) = 0

Again ( x - 8 ) is common in both terms, take it out.


==> ( x - 8 ) ( x - 2 ) = 0

Here we can take,

either ( x - 8 ) = 0 or ( x - 2 )= 0

If ( x - 8 ) = 0 ===> x = 8 and

If ( x - 2 ) = 0 ===> x = 2


So solution of given equation is,

x = 2 and

x = 8.


Hope you understood.


Regards.

praseenakos@yahoo.co.in